# Graph-Reward-SQL: Execution-Free Reinforcement Learning for Text-to-SQL via Graph Matching and Stepwise Reward

### v2

@@ Appendix I Detailed Effectiveness of GMNScore Across GRPO
N.1 CTE Rewriting Necessity Judgment
@@
N.3 CTE Rewriting Example
Table 10: Sample query and SQL responses. If it is an obvious error, we highlight the incorrect part in red.
@@
Appendix O Prompt Strategy of Text-to-SQL Task
@@
Appendix P Reasoning Prompt Strategy of Text-to-SQL Task (for Qwen2.5-Coder-7B)
@@
Appendix Q Case Study: Outcome Reward Comparison
@@
Groudtruth SELECT sum(Population) , GovernmentForm FROM country GROUP BY GovernmentForm HAVING avg(LifeExpectancy) > 72 N/A
@@
Groudtruth SELECT T1.Accelerate FROM CARS_DATA AS T1 JOIN CAR_NAMES AS T2 ON T1.Id = T2.MakeId WHERE T2.Make = ’amc hornet sportabout (sw)’; N/A
@@
Groudtruth SELECT T1.Model FROM CAR_NAMES AS T1 JOIN CARS_DATA AS T2 ON T1.MakeId = T2.Id ORDER BY T2.mpg DESC LIMIT 1; N/A
@@
Groudtruth SELECT package_option FROM TV_Channel WHERE id NOT IN (SELECT channel FROM cartoon WHERE directed_by = ’Ben Jones’) N/A

### v3

@@ Appendix I Hybrid Outcome Reward Designs
P.1 CTE Rewriting Necessity Judgment
@@
P.3 CTE Rewriting Example
@@
Appendix Q Prompt Strategy of Text-to-SQL Task
@@
Appendix R Reasoning Prompt Strategy of Text-to-SQL Task (for Qwen2.5-Coder-7B)
@@
Appendix S Case Study: Outcome Reward Comparison
@@
Groundtruth SELECT sum(Population) , GovernmentForm FROM country GROUP BY GovernmentForm HAVING avg(LifeExpectancy) > 72 N/A
@@
Groundtruth SELECT T1.Accelerate FROM CARS_DATA AS T1 JOIN CAR_NAMES AS T2 ON T1.Id = T2.MakeId WHERE T2.Make = ’amc hornet sportabout (sw)’; N/A
@@
Groundtruth SELECT T1.Model FROM CAR_NAMES AS T1 JOIN CARS_DATA AS T2 ON T1.MakeId = T2.Id ORDER BY T2.mpg DESC LIMIT 1; N/A
@@
Groundtruth SELECT package_option FROM TV_Channel WHERE id NOT IN (SELECT channel FROM cartoon WHERE directed_by = ’Ben Jones’) N/A